OpenStudy (anonymous):

prove the identity inxtanx=(cosx)/(cot^2x)

OpenStudy (anonymous):

sinx**

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

show me step by step please

OpenStudy (zzr0ck3r):

\[\sin(x)\tan(x)=\sin(x)*\frac{\sin(x)}{\cos(x)}=\frac{\sin^2(x)}{\cos(x)}=\cos(x)*\frac{\sin^2(x)}{\cos^2(x)}\\=cos(x)*\frac{1}{\frac{\cos^2(x)}{\sin^2(x)}} =\frac{\cos(x)}{\cot^2(x)}\]

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

could you help me with one more?

OpenStudy (zzr0ck3r):

sure

OpenStudy (anonymous):

(1+tanx)^2=sec^2x+2tanx

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):